### v1

# Multiplexed quantum repeaters with hot multimode alkali-noble gas memories

Abstract
We propose a non-cryogenic optical quantum memory for noble-gas nuclear spins based on the Atomic Frequency Comb (AFC) protocol. Owing to the hours-long coherence lifetimes of the noble-gas spins and the large bandwidth provided by the AFC independently of the optical depth, we estimate a time-bandwidth product of up to $9.7\times 10^{15}$ for a realistic experimental configuration. Leveraging on this long-lived multimode memory, we propose a ground-based quantum repeater that enables entanglement distribution over distances up to $2300\ \mathrm{km}$ with only $8$ elementary links. Finally, we discuss how these quantum memories can enhance rates in space-based quantum communication.
Introduction.— Optical quantum memories are a crucial building block for quantum repeaters Duan et al. (2001); Sangouard et al. (2011) enabling e.g. long-distance quantum key distribution Kimble (2008) and secure direct communication Zhang et al. (2017) or interferometric telescopes with arbitrarily long baselines Gottesman et al. (2012). A quantum memory can simultaneously act as a single-photon source Chen et al. (2006); Dideriksen et al. (2021), synchronize independent optical photons Makino et al. (2016); Davidson et al. (2023) and coherently filter them Gao et al. (2019), with applications in quantum information processing Nunn et al. (2013), including linear optical quantum computing Kok et al. (2007). In quantum sensing applications, quantum memories can be used to increase the acquisition time, for example enhancing sensitivity in NMR spectroscopy Zaiser et al. (2016) and spatial resolution in nuclear spin imaging for fundamental biology and drug discovery Ajoy et al. (2015).
The Atomic Frequency Comb (AFC) protocol Afzelius et al. (2009); de Riedmatten et al. (2008) is one of the most promising quantum memory schemes proposed to date, demonstrating high efficiency Sabooni et al. (2013) over impressive bandwidths Davidson et al. (2020) with unprecedented multi-modality Seri et al. (2019) and long storage times Ortu et al. (2022) in rare-earth ion-doped crystals at cryogenic temperatures. Atomic frequency comb based quantum memories are also promising for applications unrelated to telecommunications, including the optical detection of ultrasound, enabling precise, real-time imaging of soft biological tissues, used namely for early cancer detection McAuslan et al. (2012).
A room-temperature AFC memory for light was recently realized with alkali gas spins in cesium Main et al. (2021a) and rubidium Schofield et al. (2023) vapors, promising to extend the range of its applications by significantly reducing the complexity of experimental setups as a non-cryogenic platform, thereby increasing its robustness and scalability.
Odd isotopes of noble-gas atoms, such as ${}^{3}\mathrm{K}$ and ${}^{129}\mathrm{Xe}$, have nuclear spins exhibiting relaxation lifetimes surpassing hundreds of hours Gentile et al. (2017), corresponding to the longest-living macroscopic quantum systems presently known Katz et al. (2020a). Although optically inaccessible, they can be interfaced with light, with alkali atoms acting as mediators Katz et al. (2020b) through stochastic spin-exchange collisions O. Katz, R. Shaham and O. Firstenberg (2022).
In this Letter, we propose a non-cryogenic optical quantum memory for noble-gas nuclear spins based on the atomic frequency comb protocol. Owing to the exceptionally long coherence lifetime of the noble-gas spins and the intrinsic multi-modality of the AFC Nunn et al. (2008) for time-bin encoded qubits, our proposed scheme would enable long-lived storage and retrieval of multiple photonic modes at room temperature, with an estimated time-bandwidth product of up to $9.7\times 10^{15}$, three orders of magnitude higher than a previous proposal for this system Ji et al. (2023).
Finally, we show that a quantum repeater architecture that operates without cryogenics using these multimode quantum memories and current technology can distribute entanglement over long distances with high fidelity.
Quantum memory protocol.— We consider an ensemble of $N_{a}$ alkali-metal and $N_{b}$ noble-gas atoms in a glass cell of volume $V$ inside an optical cavity with one partially transmitting mirror, as shown in Figure 1. Initially, the alkali spins are polarized along $\hat{z}$ through optical pumping while the noble-gas spin ensemble is hyper-polarized through spin-exchange optical pumping Walker and Happer (1997) techniques.
Figure 1: (Color online) An optical quantum memory implemented on a mixture of alkali-metal (red) and noble-gas (blue) atoms contained in a spherical glass cell inside an optical cavity, with one partially transmitting mirror (the input-output coupler). A quantum signal field ^$\hat{\mathcal{E}}_{\mathrm{in}}$ interacts with the alkali atoms, which coherently couple to the noble-gas atoms through stochastic spin-exchange collisions (orange arrows).
We model the alkali-metal atoms as $\Lambda$-systems, with ground state |$\ket{g}$ and intermediate state |$\ket{s}$ both optically connected to the excited state, |$\ket{e}$, and each noble-gas atom as a two-level system with ground state $\ket{\Downarrow}$ and excited state $\ket{\Uparrow}$. A quantum light field $\hat{\mathcal{E}}$ couples the ground and excited states, while a classical control field with Rabi frequency $\Omega$ couples to the |$\ket{s}-\ket{e}$ transition.
An atomic frequency comb can be prepared in the inhomogeneously-broadened |$\ket{g}-\ket{e}$ transition via piecewise adiabatic passage techniques Rubio et al. (2018) or velocity-selective optical pumping with several narrowband frequency light modes Main et al. (2021b) or an optical frequency comb from a mode-locked laser Aumiler et al. (2005), where the peak separation is given by the laser repetition rate. We suppose that an atomic frequency comb with peak width (FWHM) $\gamma$, separation $\Delta$ and overall width $\Gamma$ is prepared by either of these methods, and denote its finesse by $F=\Delta/\gamma$.
The evolution of the (slowly-varying) cavity field $\hat{\mathcal{E}}$ is then governed by the Heisenberg-Langevin equation Afzelius and Simon (2010)
[…]
where $\kappa$ is the cavity decay rate, $g_{0}$ is the atom-cavity coupling constant, Weierstrass-$\wp$ is the dipole moment for the |$\ket{g}-\ket{e}$ transition, $n(\delta)$ is the atomic spectral distribution and $\hat{\sigma}_{ge}$ is the atomic polarization at detuning $\delta$, obeying
$\partial_{t}\hat{\sigma}_{ge}=-(i\delta+\gamma_{p})\hat{\sigma}_{ge}+i\wp\hat{\mathcal{E}},$ (2)
where $\gamma_{p}$ is the homogeneous linewidth of the optical transition. In the fast cavity limit, the incoming and outgoing cavity fields obey the input-output relation
^$\hat{\mathcal{E}}_{\mathrm{out}}-\hat{\mathcal{E}}_{\mathrm{in}}=\sqrt{2\kappa}\hat{\mathcal{E}}.$ (3)
Accordingly, we consider the dynamics of the collective spin operators for the alkali and noble-gas ensembles,
$\ \hat{\mathcal{S}}(\mathbf{r},t)=\frac{\hat{\sigma}_{gs}(\mathbf{r},t)}{\sqrt{p_{a}n_{a}}},\qquad\hat{\mathcal{K}}(\mathbf{r},t)=\frac{\hat{\sigma}_{\Downarrow\Uparrow}(\mathbf{r},t)}{\sqrt{p_{b}n_{b}}},$ (4)
where $\hat{\sigma}_{\alpha\beta}$ are continuous atomic coherences Sangouard et al. (2007), $p_{a},p_{b}$ denote the polarization degree and $n_{a},n_{b}$ the atomic density of each ensemble Katz et al. (2022), along with the optical dipole
[…]
In a frame co-rotating with the quantum signal field ^$\hat{\mathcal{E}}_{\mathrm{in}}$, assuming that the excited state population remains negligible, for highly polarized ensembles $p_{a},p_{b}\approx 1$, we obtain the equations of motion for these bosonic operators Katz et al. (2022),
$\partial_{t}\hat{\mathcal{P}}$ $=-(\gamma_{p}+i\bar{\delta})\hat{\mathcal{P}}+i\Omega\hat{\mathcal{S}}+iG\hat{\mathcal{E}}+\hat{f}_{\mathcal{P}},$ (6)
$\partial_{t}\hat{\mathcal{S}}$ $=-(\gamma_{s}+i\delta_{s}-D_{a}\nabla^{2})\hat{\mathcal{S}}+i\Omega^{*}\hat{\mathcal{P}}-iJ\hat{\mathcal{K}}+\hat{f}_{\mathcal{S}},$ (7)
$\partial_{t}\hat{\mathcal{K}}$ $=-(\gamma_{k}+i\delta_{k}-D_{b}\nabla^{2})\hat{\mathcal{K}}-iJ\hat{\mathcal{S}}+\hat{f}_{\mathcal{K}},$ (8)
where $\bar{\delta}$ is the average optical detuning, $\delta_{s}$ and $\delta_{k}$ are the two-photon detunings from the alkali and noble-gas spin resonances, $\gamma_{p}\gg\gamma_{s}\gg\gamma_{k}$ are the decoherence rates for the optical dipole, alkali and noble-gas spin coherences, $D_{a}$ and $D_{b}$ are the diffusion coefficients for the alkali and noble-gas species, $G=\sqrt{p_{a}n_{a}}g_{0}\wp$ is the collective atom-field coupling frequency and $J=\zeta\sqrt{p_{a}p_{b}n_{a}n_{b}/4}$ is the coherent spin-exchange rate, where $\zeta$ is a constant O. Katz, R. Shaham and O. Firstenberg (2022).
Here, the quantum noise processes $\hat{f}_{\mathcal{P}}$, $\hat{f}_{\mathcal{S}}$ and $\hat{f}_{\mathcal{K}}$ ensure the preservation of the commutation relations of the corresponding operators by introducing vacuum noise Shaham et al. (2020).
At first, with the control field turned off ($\Omega=0$), the quantum light field is absorbed by the alkali ensemble and it can then be shown from equations (1)-(3) that
^$\hat{\mathcal{E}}_{\mathrm{out}}=\frac{\kappa-Z}{\kappa+Z}\ \hat{\mathcal{E}}_{\mathrm{in}}.$ (9)
where $Z=N_{a}g_{0}^{2}\wp^{2}/\Gamma$ is the absorption rate of the cavity field by the atomic ensemble. If the cavity absorption and decay rates match, $\kappa=Z$, the input field is fully absorbed, ^$\hat{\mathcal{E}}_{\mathrm{out}}=0$ Afzelius and Simon (2010). Notably, this impedance matching condition corresponds to unit optical cooperativity, $C=N_{a}G^{2}/\kappa\Gamma=1$, a modest experimental constraint.
In the second stage of the protocol, a chirped adiabatic pulse is used to transfer the optical coherence onto the alkali spin-wave $\hat{\mathcal{S}}$. A hyperbolic-square-shaped (HSH) pulse Tian et al. (2011) with duration $T$ can be employed to achieve efficient broadband population transfer Businger et al. (2020), such that
$\langle\hat{\mathcal{S}}^{\dagger}\hat{\mathcal{S}}\rangle_{T}=\left[1-\mathrm{exp}\left(-\frac{\pi^{2}T\Omega^{2}}{\Gamma}\right)\right]\langle\hat{\mathcal{P}}^{\dagger}\hat{\mathcal{P}}\rangle_{0}.$ (10)
Then, the collective alkali excitation is coherently transferred to the noble-gas ensemble through stochastic spin-exchange collisions O. Katz, R. Shaham and O. Firstenberg (2022) after $T^{\prime}\approx(\pi J-\gamma_{s})/(2J^{2})$,
$\langle\hat{\mathcal{K}}^{\dagger}\hat{\mathcal{K}}\rangle_{T^{\prime}}=\mathrm{exp}\left(-\frac{\pi\gamma_{s}}{2J}\right)\langle\hat{\mathcal{S}}^{\dagger}\hat{\mathcal{S}}\rangle_{T}.$ (11)
A strong magnetic field ($\delta_{k}\gg J$) is subsequently applied to effectively decouple the ensembles. In the dark, the noble-gas spins function as a very long-lived quantum memory, whose storage time is bounded only by $1/\gamma_{k}$, potentially exceeding hundreds of hours Katz et al. (2020a, 2022). Finally, the magnetic field is switched off and the storage sequence $\hat{\mathcal{P}}\to\hat{\mathcal{S}}\to\hat{\mathcal{K}}$ is reversed. At a time $2\pi/\Delta$ after the reversal is concluded, the periodic structure of the comb leads to a coherent re-emission of the stored pulse,
^$\hat{\mathcal{E}}_{\mathrm{out}}(t)=-\sqrt{\eta_{m}}\ \hat{\mathcal{E}}_{\mathrm{in}}\left(t-2T^{\prime}-2T-\frac{2\pi}{\Delta}\right),$ (12)
with the memory efficiency $\eta_{m}$ is approximately given by
$\eta_{m}=\left[1-\mathrm{exp}\left(-\frac{\pi^{2}T\Omega^{2}}{\Gamma}\right)\right]^{2}\mathrm{exp}\left(-\frac{\pi\gamma_{s}}{J}\right)\mathrm{sinc}^{2}\left(\frac{\pi}{F}\right)$ (13)
for an atomic frequency comb with square peaks Bonarota et al. (2010).
Figure 2: (Color online) (a) Expectation values of the alkali $\langle\hat{\mathcal{S}}^{\dagger}\hat{\mathcal{S}}\rangle$ and noble-gas $\langle\hat{\mathcal{K}}^{\dagger}\hat{\mathcal{K}}\rangle$ (b) spin populations throughout the memory protocol, with $D_{a}=0.35\ \mathrm{cm}^{2}\ \mathrm{Hz}$, $D_{b}=0.70\ \mathrm{cm}^{2}\ \mathrm{Hz}$, $\delta_{k}=50\ \mathrm{kHz}$ and the parameters described in the main text. (c) Storage and retrieval of a sequence of Gaussian pulses with duration $\tau=30\ \mathrm{ns}$. (d) Memory efficiency estimated from equation (13) (dashed line) and from the numerical simulation of equations (7)-(8) (solid line) as a function of $J/\gamma_{s}$.
Implementation.— We consider a ${}^{3}\mathrm{He-K}$ mixture confined in an uncoated spherical glass cell with radius $1\ \mathrm{cm}$ at $215^{\circ}\mathrm{C}$, in a high-pressure configuration proposed in Ref. O. Katz, R. Shaham and O. Firstenberg (2022). We assume a ${}^{39}\mathrm{K}$ density of $n_{a}=3\times 10^{14}\ \mathrm{cm}^{-3}$ (at vapor pressure), a ${}^{3}\mathrm{He}$ density of $n_{b}=2\times 10^{20}\ \mathrm{cm}^{-3}$, obtaining a coherent spin-exchange rate of $J=1000\ \mathrm{Hz}$ for $p_{a}=0.95$ and $p_{b}=0.75$ Chen et al. (2014). A buffer gas ($\mathrm{N}_{2}$) is introduced for quenching, at a pressure of $30\ \mathrm{Torr}$, broadening the potassium $\mathrm{D}_{1}$ linewidth to $27\ \mathrm{GHz}$. In practice, the AFC peak width $\gamma$ is limited by the optical coherence lifetime, bounded by twice the natural linewidth of the transition, $\gamma_{p}=5.96\times 2\pi\ \mathrm{MHz}$. An intrinsic AFC efficiency of $95\%$ can be achieved for a finesse $F=8$, corresponding to a peak separation $\Delta=96\ \mathrm{MHz}$. Assuming the frequency comb bandwidth matches the broadened transition linewidth, we obtain an upper bound on the multimode capacity of $M=2\Gamma/5\Delta=112$ Ortu et al. (2022). In this configuration, the alkali relaxation rate is estimated to be $\gamma_{s}=17.5\ \mathrm{Hz}$ O. Katz, R. Shaham and O. Firstenberg (2022). A population transfer efficiency of $98\%$ can be attained using HSH pulses with $\pi^{2}T\Omega^{2}/\Gamma=4$ Ortu et al. (2022). We therefore estimate a memory efficiency of $\eta_{m}=88\%$ using equation (13), mainly limited by the spin-exchange efficiency.
In Figure 2, we represent the expectation values of the collective spins of the alkali-metal and noble-gas ensembles throughout the memory protocol, obtained through the numerical solution of equations (7)-(8) with destructive (Dirichlet) and non-destructive (von Neumman) boundary conditions for the alkali and noble-gas spins, respectively Shaham et al. (2020), representing the interaction with the surface of the glass cell. We estimate a memory efficiency of $79\%$ from our numerical solution, compared to an analytical prediction of $88\%$ from equation (13), which is accurate only for anti-relaxation coated cells Katz et al. (2022). In uncoated cells, the collective spin operators are subject to distinct boundary conditions so that their spatial modes only partially overlap; this mode mismatch induced by diffusion further reduces the spin-exchange efficiency.
Fiber-based quantum repeater.— We consider a quantum repeater with photon pair sources and our proposed temporally multimode memories, using alkali and noble-gas atoms confined in a vapor cell. A photon pair source can also be realized using warm vapor cells Park et al. (2019); Kim et al. (2022) or parametric down-conversion in nanophotonic chips Guo et al. (2016).
As a first step, to generate entanglement between two nodes $A$ and $B$, the pair sources at both locations are coherently and simultaneously excited, such that each of them has a probability $p/2\ll 1$ of creating a pair. Then, for each pair, one mode is stored in the quantum memory while the other is coupled into an optical fiber and combined in a beam splitter Sangouard et al. (2011); Simon et al. (2007). Hence, the detection of a photon after the beam splitter heralds the storage of a single photon, projecting the system into an entangled state with a single delocalized atomic excitation Duan et al. (2001),
[…]
In effect, using multimode memories, entanglement creation can be attempted multiple times in each clock interval $L_{0}/c$, thus increasing the success probability $P$ to $1-(1-P)^{N}\approx NP$, for $N$ distinct temporal modes Simon et al. (2007).
Figure 3: (Color online) Entanglement distribution rates for a fiber-based quantum repeater as a function of distance for 4 and 8 links employing our proposed scheme (solid lines), the single-photon protocol of Ji et al. Ji et al. (2023) (short-dashed lines), the standard DLCZ protocol Duan et al. (2001) (long-dashed lines) and direct transmission with a source rate of $10\ \mathrm{GHz}$ (dashed line), using the parameters described in the main text for all curves.
Finally, the entanglement can be teleported to a distant node by successive entanglement swapping. The total time required for entanglement distribution is Sangouard et al. (2011)
$T_{\mathrm{tot}}=\left(\frac{L_{0}}{c}+\frac{\pi}{2J}\right)\frac{3^{n+1}}{Np\eta_{c}\eta_{d}\eta_{t}\eta^{n+2}}\ \prod_{k=1}^{n}\left(2^{k}(2^{k}-1)\eta\right),$ (15)
where $L_{0}$ is the elementary link length, $n=\log_{2}M$ is the nesting level of a repeater with $M$ nodes, $\eta_{d}$ is the detection efficiency, $\eta_{c}$ is the telecommunication frequency conversion efficiency, $\eta=\eta_{m}\eta_{d}$, $\eta_{t}=\mathrm{e}^{-L_{0}/(2L_{\mathrm{att}})}$ is the transmission efficiency, $L_{\mathrm{att}}=22\ \mathrm{km}$ is the attenuation length and $c=2.0\times 10^{8}\ \mathrm{m/s}$ is the speed of light in the optical fiber. We consider using silicon single-photon avalanche diodes with detection efficiency $\eta_{d}=0.75$ Fang et al. (2020) and lithium niobate waveguide chips for frequency conversion Zheng et al. (2020) assuming $\eta_{c}=0.8$, as proposed in Ref. Ji et al. (2023).
A wafer of microfabricated vapor cells could provide for hundreds of independent memories in each location, speeding up entanglement distribution. An optical quantum memory in a microfabricated vapor cell has been recently demonstrated Mottola et al. (2023), showing that significant parallelization can be achieved with near-term technology.
In Figure 3, we plot the entanglement distribution rates for our proposed quantum repeater architecture with $100$ memories at each node, comparing it with the single-photon proposal of Ref. Ji et al. (2023) and the well-known DLCZ Duan et al. (2001) protocol for atomic ensembles. Although the long storage and retrieval times $\pi/2J\gtrsim L_{0}/c$ due to the slow interface between the alkali and noble-gas spins considerably reduce the entanglement distribution rates, the temporal multiplexing enabled by the atomic frequency comb can easily compensate for this limitation, as Figure 3 shows, clearly surpassing the original DLCZ protocol Duan et al. (2001) rates for a fast memory. We find that our proposed protocol outperforms direct transmission for distances over $507\ \mathrm{km}$ and offers higher rates than the single-photon architecture proposed in Ref. Ji et al. (2023) for the hybrid alkali noble-gas memories we consider. A $4$-link repeater is optimal for distances under $1400\ \mathrm{km}$, whereas an $8$-link repeater can achieve better rates for longer distances. Ultimately, the achievable distance is limited by the memory storage times, since quantum repeater protocols require it to be comparable to $T_{\rm tot}$ so that post-selection of the entangled pair is possible Sangouard et al. (2011). In the configuration we considered, the collective ${}^{3}\mathrm{He}$ nuclear spin can remain coherent for $100$ hours O. Katz, R. Shaham and O. Firstenberg (2022), enabling entanglement distribution for distances up to $2280\ \mathrm{km}$. In practice, due to fiber losses this is believed to be close to the limit for terrestrial quantum networks without error correction Simon (2017), which requires considerably more complex resources.
Conclusion.— In this Letter, we have proposed a non-cryogenic, multimode quantum memory based on atomic frequency combs in alkali-noble gas ensembles. With our proposed memory scheme, a quantum repeater that outperforms direct transmission could be demonstrated with only $4$ links, operating at room temperature. Overcoming this barrier would constitute an important technological milestone for ground-based quantum repeaters Simon (2017)
A solution to overcome fiber losses and achieve even longer distances is to use satellite-borne photon sources Yin et al. (2017a) or quantum memories Gündoğan et al. (2021), since photon absorption and decoherence in empty space are negligible. An entangled photon pair has been distributed to locations separated by $1200\ \mathrm{km}$ in a space-to-ground two-downlink channel Yin et al. (2017b), with losses as low as $64\ \mathrm{dB}$. However, satellite links are plagued by intermittency due to weather and satellite-orbit shadow time, which considerably reduce entanglement distribution rates. Hence, very long-lived quantum memories such as we propose here can alleviate these issues, storing entangled pairs when the satellite link is available Simon (2017), further increasing rates due to the temporal multiplexing enabled by the AFC. In this way, our proposed memory scheme could be incorporated in hybrid or satellite-based repeater schemes Gündoğan et al. (2021), thereby contributing towards the longstanding goal of building a quantum network spanning truly global distances.
Acknowledgements.— The authors acknowledge Fundação para a Ciência e a Tecnologia (FCT-Portugal) through national funds and when applicable co-funding by EU funds under the Projects No. UIDB/50008/2020, PTDC/FIS-OUT/3882/2020, and Contract Nos. CEECIND/00401/2018 and CEECIND/CP1653/CT0002.